The average age of bank managers is 35 years. Assume the variable is normally distributed. If the standard deviation is 5 years, find the probability that the age of a randomly selected bank manager will be in the range between 25 and 40 years old.


1
Expert's answer
2021-12-02T11:02:27-0500

"P(25<X<40)=P(\frac{25-35}{5}<Z<\frac{40-35}{5})"

"=P(-2<Z<1)"

"=P(Z<1)-P(Z<-2)"

From normal standard distribution table;

"P(25<X<40)=0.8413-0.0227"

"P(25<X<40)=0.8186"